OpenStudy (nickszuhy):

What is the equation of the directrix of the parabola given by the equation (y − 3)^2 = 8(x − 5)?

OpenStudy (mathmale):

When we're discussing the focus and directrix of a horizontal parabola whose vertex is at the origin (0,0), the appropriate equation is 4px=y^2. If, however, the vertex is at (h,k), the appropriate equation is 4p(x-h)=(y-k)^2. Compare this standard form with your specific equation of a parabola. What are your p, your h, your k?